Grade 9Mechanics

a machine gun of mass 12 kg fires bullets of mass 25g at the rate of 4 bullets/sec......with a velocity 500m/s...........what force must be applied to the gun to hold it in position.....??????????????????????PLS EXPLAIN!!!

Shivam Bhagat
14 Years agoGrade 9

1 Answer

Rahul Kumar
14 Years ago

First of all you should know that all Physics Qs are based on ideal conditions,so the momentum(linear) always remains conserved....

ie m1v1=m2v2

in this case

GUN                BULLET

m1=12kg        m2=25gm=>0.025kg=>(0.025*4)kg per sec

v1=x(let)m/s  v2=500m/s

since momentum is conserved

12*x=0.025*4*500

x=(2.5*4*5)/12

x=4.16667m/s

F=change in momentum,so

12*4.1667-0="GIVES YOU THE CORRECT ANSWER!!"